# Pre-trained Sequential Recommendation Framework via Popularity Dynamics

### v2

@@ 5.1. Datasets and Preprocessing
  • Amazon (Ni et al., 2019) is a series of product ratings datasets obtained from Amazon.com, split by product domain. We consider the Office and Tool product domains in our study.
  • Douban (Song et al., 2019) consists of three datasets across different domains, collected from Douban.com, a Chinese review website. We work with the Movie and Music datasets.
  • Epinions (Tang et al., 2012a, b) is a dataset crawled from product review site Epinions. We utilize the ratings dataset for our study.
We present dataset statistics in Table 1. We compute the density as the ratio of the number of interactions to the number of users times the number of items. Both datasets from Douban (i.e., movie and music) are the densest and have no auxiliary information available, while the Amazon review datasets (i.e., office and tool) are the sparsest.
For fair evaluation, we follow the same preprocessing procedure as previous works (Kang and McAuley, 2018; Sun et al., 2019; Rendle et al., 2010; Tang and Wang, 2018), i.e., we binarize the explicit ratings to implicit feedback. In addition, for each user, we sort interactions by their timestamp and take these chronological sequences as input, using the second most recent action for validation, and the most recent action for testing.
@@
MostPop
0.450
0.272
0.459
0.274
0.586
0.361
0.519
0.327
0.438
0.296
BPR (Rendle et al., 2009)
0.457
0.289
0.363
0.216
0.747
0.477
0.646
0.434
0.568
0.397
NCF (He et al., 2017)
0.446
0.266
0.388
0.239
0.784
0.505
0.652
0.437
0.570
0.396
NGCF (Wang et al., 2019)
0.460
0.291
0.421
0.256
0.788
0.510
0.659
0.441
0.571
0.393
LightGCN (He et al., 2020)
0.465
0.293
0.463
0.275
0.793
0.512
0.665
0.447
0.575
0.396
@@
Caser (Tang and Wang, 2018)
0.512
0.334
0.496
0.297
0.891
0.701
0.796
0.576
0.674
0.475
MARank (Yu et al., 2019)
0.488
0.312
0.468
0.282
0.888
0.683
0.739
0.501
0.628
0.436
SasRec (Kang and McAuley, 2018)
0.539
0.354
0.536
0.337
0.918
0.749
0.816
0.599
$0.705^{*}$
0.501
BERT4Rec (Sun et al., 2019)
$0.541^{*}$
$0.358^{*}$
$0.544^{*}$
$0.350^{*}$
0.900
0.728
$0.816^{*}$
$0.602^{*}$
0.702
$0.512^{*}$
TiSasRec (Li et al., 2020)
0.531
0.349
0.539
0.341
$0.918^{*}$
$0.752^{*}$
0.809
0.523
0.701
0.499
PrepRec
0.536
0.344
0.551
0.339
0.908
0.738
0.782
0.573
0.795
0.580
PrepRec $\Delta$
-0.9%
-3.9%
+1.2%
-3.1%
-1.1%
-1.8%
-1.9%
-4.8%
+12.7%
+13.3%
@@
Table 3. Regular sequential recommendation results. We embolden the best results and mark the best baseline results with $*{}^{\prime}*^{\prime}$. Interp represents the interpolation results between PrepRec and BERT4Rec. PrepRec $\Delta$ denotes the performance difference between PrepRec and the best results among the selected baselines, similar for Interp $\Delta$. Despite not focusing on the performance of regular sequential recommendation, PrepRec achieves comparable performance to the state-of-the-art sequential recommenders, with only on average 0.2% worse than the best performing sequential recommenders in R@10 while having only a fraction of the model size (Table 2). After a simple post-hoc interpolation, we outperform the state-of-the-art sequential recommenders by 13.8% in R@10 on average.
5.2. Baselines
To assess the effectiveness of our proposed method, we compare against the following set of baselines. These baselines include classic general recommendation models and state-of-the-art sequential recommendation models.
  • MostPop is a non-personalized baseline that recommends the most popular items in the training set.
  • BPR (Rendle et al., 2009) learns latent representations for users and items through matrix factorization.
  • NCF (He et al., 2017) learns user and item embeddings through a multi-layer perceptron instead of matrix factorization.
  • NGCF (Wang et al., 2019) uses a bipartite graph to model user-item interactions through a graph convolutional network (Kipf and Welling, 2016).
  • LightGCN (He et al., 2020) is a simplified version of NGCF that removes the feature transformation and nonlinear activation.
  • Caser (Tang and Wang, 2018) learns convolutional filters on user interaction sequences over the sequence order and latent dimensions.
  • MArank (Yu et al., 2019) captures individual and union-level item dependency through multi-order attention.
  • SasRec (Kang and McAuley, 2018) uses self-attention to model user embeddings through its sequence of latent item representations.
  • BERT4Rec (Sun et al., 2019) improves on SasRec by using a bi-directional transformer to capture interactions from both sides.
  • TiSasRec (Li et al., 2020) additionally embeds the raw time interval between consecutive item interactions in the sequence.
To the best of our knowledge, we are the first to tackle the zero-shot sequential recommendation without any auxiliary information. (Hou et al., 2022, 2023; Wang et al., 2023b) are designed for few-shot adaptation and uses auxiliary information. (Ding et al., 2021) is capable of zero-shot recommendation but requires textual descriptions of items, which we do not consider in the experiments. Therefore, we do not include the existing pre-trained sequential recommenders as baselines.
5.3. Experimental Setup
@@
Following previous works (Koren, 2008; He et al., 2017; Kang and McAuley, 2018; Sun et al., 2019), we adopt the leave-one-out evaluation method: for each user, we pair its test item with 100 unobserved items according to the user’s interaction history. Then we rank the test item for the user among the 101 total items. We use two standard evaluation metrics for top-$k$ recommendation: Recall@k (R@k) and Normalized Discounted Cumulative Gain@N (N@k). R@k measures the proportion of test items that appear in the top-$k$ recommendations, and N@k accounts for the ranking position of the test interaction. Our model explicitly utilize popularity information. Therefore, in addition to randomly sampled unobserved items (negatives), we also present results where we sample the negatives based on their popularities, i.e., popular items have higher probabilities of being sampled as negatives, to show that PrepRec can learn discriminative item representations. We report the average of R@k and N@k over all the test interactions.
@@
We use publicly available Pytorch implementations for the baselines. For fair evaluation, we set dimension size $d$ to 50, max sequence length $L$ to $200$, and batch size to $128$ in all models. We use three layers (one base and two additional) in NCF and NGCF, and two attention layers for the sequential models (SasRec, BERT4Rec, TiSasRec, and ours). We use an Adam optimizer and tune the learning rate in the range $\{10^{-4},10^{-3},10^{-2}\}$ and set the weight decay to $10^{-5}$. We use the dropout regularization rate of $0.3$ for all models. We set $\gamma=0.5$ in Equation 1, whose reason we will discuss the reason in  § 5.8.1. We define the coarse and fine period to be 10 and 2 days respectively, and we fix the window size to be $m=12$ and $n=4$ for all datasets (§ 4.1.2). We train PrepRec for a maximum of 80 epochs. All experiments are conducted on a Tesla V100 using PyTorch. We repeat each experiment 5 times with different random seeds and report the average performance.
@@
Metric
R@10
N@10
R@10
N@10
R@10
N@10
MostPop
0.197
0.139
0.099
0.046
0.163
0.110
SasRec (Kang and McAuley, 2018)
0.749
0.519
0.453
0.291
0.658
0.442
BERT4Rec (Sun et al., 2019)
0.747
0.519
0.461 0.299
0.655
0.456
PrepRec
0.739
0.523
0.443
0.280
0.762 0.551
PrepRec $\Delta$
-1.3%
+0.7%
-2.2%
-6.3%
+15.8%
+17.2%
@@
Table 4. Regular sequential recommendation results with popularity-based negative sampling. We present additional in-domain recommendation results with unobserved (negative) items sampled based on their popularity (Sun et al., 2019) to show that PrepRec can learn discriminative universal item and sequence representations despite depending only on popularity statistics.
[…]
Metric
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
Reference: Regular Sequential Recommendation
MostPop
0.450
0.272
0.459
0.274
0.586
0.361
0.519
0.327
0.438
0.296
PrepRec
0.536
0.344
0.551
0.339
0.908
0.738
0.782
0.573
0.795
0.580
[…]
Office $\rightarrow$
0.540 0.326
0.838
0.624
0.755
0.542
0.724
0.512
Tool $\rightarrow$
0.543 0.332
0.881 0.659
0.749
0.536
0.717
0.510
[…]
0.520
0.320
0.508
0.302
0.811 0.600 0.751 0.537
[…]
0.503
0.310
0.496
0.312
0.836
0.636
0.739
0.518
[…]
0.517
0.317
0.470
0.302
0.872
0.656
0.774
0.517
[…]
5.4. Regular Sequential Recommendation (RQ1)
We show the performance of PrepRec against the selected baselines. We want to emphasize that the purpose of this paper is not to outperform the state-of-the-art sequential recommenders but to set a basline for further pre-trained sequential recommenders. We include the results here to investigate how well the popularity dynamics can items and sequences in the sequential recommendation. PrepRec achieves competitive performance- within 2% in R@10 and 5% in N@10, - with the state-of-the-art sequential recommenders. On Epinions, PrepRec even outperforms all the sequential recommenders by 7.3%. This is particularly impressive given that PrepRec has significantly fewer model parameters (Table 2).
Since PrepRec explicitly models popularity information and the MostPop demonstrates competitive performance compared to the remaining baselines, we conduct an additional experiment (Table 4) where we sample the unobserved (negative) items based on their popularity (Sun et al., 2019). As shown in Table 4, MostPop’s performance dropped significantly, while PrepRec shows even more competitive performance on some datasets (e.g., Music and Epinions). This suggests even though PrepRec utilizes non-personalized popularity information, it can still learn discriminative item and sequence representations.
PrepRec learns item representations through popularity dynamics, which is conceptually different from learning representations specific to each item ID. Therefore, we propose a simple post-hoc interpolation to investigate if modeling popularity dynamics can boost the performance of state-of-the-art sequential recommenders. We interpolate the scores from PrepRec with the scores from BERT4Rec as follows: $\hat{y}_{intp}(v^{t}_{j}|\mathcal{S}_{u})=\alpha*\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})+(1-\alpha)*\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$, where $\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})$ and $\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ are the scores from PrepRec (Equation 10) and BERT4Rec, respectively. We set $\alpha=0.5$ for all datasets. After interpolation, there is a significant performance boost by up to 34.9% in N@10. This suggests that our model effectively captures dynamic item properties that are missing with the static item embeddings of SOTA sequential models. Gains are largest in the medium and low-density datasets (Epinions, Amazon), indicating that our model is particularly complimentary to existing methods in sparse datasets where item embeddings are less informative.
5.5. Zero-shot Transfer(RQ2)
We follow the zero-shot inference setting introduced in § 4.3 and report the results in Table 5. We also include the results of the regular sequential recommendation for reference. In the zero-shot setting, after training on source datasets, PrepRec shows minimal performance reduction in the target datasets (i.e., 6% maximum and 2% average reduction in R@10). The best zero-shot transfer results from PrepRec only fall short against the selected sequential recommendation baselines (Table 3) by up to 4% and even outperform (by up to 6.5%) them on the target datasets (e.g., Epinions and Office). We found that our models trained on Movie and tools show the highest generalizability, even outperforming the target-trained models on Music (0.811 vs. 0.782 on R@10). We conjecture that this is because Movie is the largest dataset in terms of the number of interactions. Overall, these results show PrepRec ’s effectiveness in zero-shot transfer without any training on interaction data or side information. In addition, this experiment also demonstrates that the popularity dynamics-based item and sequence representations are generalizable across domains and applications.
5.6. Ablation Study (RQ3)
@@
Metric
R@10
N@10
R@10
N@10
R@10
N@10
PrepRec 0.782 0.573
0.536
0.344 0.795 0.580
w/o Relative Time ${\bm{T}}$ (§ 4.1.3)
0.734
0.514
0.541
0.334
0.782
0.562
w/o Positional ${\bm{P}}$ (§ 4.1.4)
0.765
0.544
0.530
0.332
0.772
0.554
Sinuoisal Popularity Encoding
0.779
0.570
0.529
0.340
0.772
0.561
@@
Here, we assess the importance of different components crucial to PrepRec , i.e., relative time encoding (§ 4.1.3), positional encoding (§ 4.1.4), and popularity encoder $E_{p}$ (§ 4.1.1). We find that removing relative time encoding ${\bm{T}}$ results in the largest performance drop on both the Music and Office datasets. This suggests that the relative time encoding is crucial for effectively capturing the popularity dynamics. Removing positional encoding ${\bm{P}}$ results in a maximum of 2.2% drop in R@10 on the Office dataset, indicating positional encoding is important for capturing sequential information. Finally, changing $E_{p}$ to the non-linear sunusoidal encoding shows worse performance on all datasets, meaning that the linear encoding is more suitable for capturing the popularity dynamics.
5.7. Qualitative Analysis(RQ4)
Figure 2. Recommendation results for different item popularity groups. Item Popularity Group 1 represents the least popular items, and Group 5 represents the most popular items. PrepRec achieves better performance on long-tail items, while the sequential recommenders show stronger performance on popular items.
Here we present a detailed analysis regarding the performance of PrepRec . We separate test items into equally sized groups based on their popularity in the training set. We then compute the average R@10 and N@10 for each group. We present the results in Figure 2. We find that PrepRec achieves better performance on item group with the least interactions, i.e., long-tail items, while the SasRec and BERT4Rec show stronger performance on popular items. Long-tail item recommendation is a particularly challenging task explored by many previous works (Sankar et al., 2021) and requires recommenders able to learn high-quality representations with just a few interactions. This corresponds to our observation that PrepRec is more robust to data sparsity and can learn discriminative item and sequence representations (§ 5.4), showing that long-tail item recommendation can benefit from PrepRec ’s popularity dynamics-based item representations.
5.8. Further Analysis
@@
Metric
R@10
N@10
R@10
N@10
R@10
N@10
$\gamma=0.5^{*}$ (weighted -pop)
0.782 0.573 0.536 0.344 0.795 0.580
$\gamma=0$ (Curr-pop)
0.749
0.542
0.512
0.328
0.689
0.496
$\gamma=1$ (cumul-pop)
0.695
0.452
0.530
0.330
0.733
0.505
Table 7. Recommendation results for varying the discounting factor $\gamma$ in § 4.1.2. $\gamma=0.5$ is the default setting, denoted by $*{}^{\prime}*^{\prime}$. We find that $\gamma=0.5$ generally outperforms the other two settings

### v3

@@ 5.1. Datasets and Preprocessing
Amazon (Ni et al., 2019) is a series of product ratings datasets obtained from Amazon.com, split by product categories. We consider the Office and Tool product domains in our study. Douban (Song et al., 2019) consists of three datasets across different domains, collected from Douban.com, a Chinese review website. We work with the Movie and Music datasets. Epinions (Tang et al., 2012a, b) is a dataset crawled from product review site Epinions. We utilize the ratings dataset for our study.
We present dataset statistics in Table 1. We compute the density as the ratio of the number of interactions to the number of users times the number of items. Douban datasets (i.e., movie and music) are the densest and have no auxiliary information available, while the Amazon review datasets (i.e., office and tool) are the sparsest.
For fair evaluation, we follow the same preprocessing procedure as previous works (Kang and McAuley, 2018; Sun et al., 2019), i.e., we binarize the explicit ratings to implicit feedback. In addition, for each user, we sort interactions by their timestamp and use the second most recent action for validation, the most recent action for testing, and the rest for training.
5.2. Baselines and Experimental Setup
Baselines: Our baselines (supplementary materials contain detailed descriptions) include classic general recommendation models (e.g., MostPop, BPR (Rendle et al., 2009), NCF (He et al., 2017), LightGCN (He et al., 2020)) and state-of-the-art sequential recommendation models (e.g., Caser (Tang and Wang, 2018), SasRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), TiSasRec (Li et al., 2020), CL4SRec (Xie et al., 2022)).
@@ 5.2. Baselines and Experimental Setup
MostPop 0.450 0.272 0.459 0.274 0.586 0.361 0.519 0.327 0.438 0.296
BPR (Rendle et al., 2009) 0.457 0.289 0.363 0.216 0.747 0.477 0.646 0.434 0.568 0.397
NCF (He et al., 2017) 0.446 0.266 0.388 0.239 0.784 0.505 0.652 0.437 0.570 0.396
LightGCN (He et al., 2020) 0.465 0.293 0.463 0.275 0.793 0.512 0.665 0.447 0.575 0.396
@@
Caser (Tang and Wang, 2018) 0.512 0.334 0.496 0.297 0.891 0.701 0.796 0.576 0.674 0.475
SasRec (Kang and McAuley, 2018) 0.539 0.354 0.536 0.337 0.918 0.749 0.816 0.599 $0.705^{*}$ 0.501
BERT4Rec (Sun et al., 2019) 0.541 0.358 0.544 0.350 0.900 0.728 $0.816^{*}$ $0.602^{*}$ 0.702 $0.512^{*}$
TiSasRec (Li et al., 2020) 0.531 0.349 0.539 0.341 $0.918^{*}$ $0.752^{*}$ 0.809 0.523 0.701 0.499
CL4SRec (Xie et al., 2022) $0.550^{*}$ $0.358^{*}$ $0.548^{*}$ $0.352^{*}$ 0.899 0.725 0.813 0.597 0.662 0.481
PrepRec 0.536 0.344 0.551 0.339 0.908 0.738 0.782 0.573 0.795 0.580
PrepRec $\Delta$2.5% -3.9% +1.2% -3.6% -1.1% -1.8% -1.9% -4.8% +12.7% +13.3%
@@
Table 3. Regular sequential recommendation results, RQ2, (§ 5.4.1). We make bold the best results and mark the best baseline results with ${}^{\prime}*^{\prime}$. Interp represents the interpolation results between PrepRec and BERT4Rec. PrepRec $\Delta$ denotes the performance difference between PrepRec and the best results among the selected baselines, similar for Interp $\Delta$. PrepRec achieves comparable performance to the state-of-the-art sequential recommenders, with only on average 0.2% worse than the best performing sequential recommenders in R@10 while having only a fraction of the model size (Table 5). After a simple post-hoc interpolation, we outperform the state-of-the-art sequential recommenders by 11.8% in R@10 on average.
@@
Following previous works (Koren, 2008; He et al., 2017; Kang and McAuley, 2018; Sun et al., 2019), we adopt the leave-one-out evaluation method: for each user, we pair the test item with 100 unobserved items according to the user’s interaction history. Then we rank the test item for the user among the 101 total items. We use two standard evaluation metrics for top-$k$ recommendation: Recall@k (R@k) and Normalized Discounted Cumulative Gain@k (N@k). Our model explicitly utilizes popularity information. Therefore, we also present results where we sample the negatives based on their popularities, i.e., popular items have higher probabilities of being sampled as negatives. We report the average of R@k and N@k over all the test interactions.
@@
We use publicly available implementations for the baselines. For fair evaluation, we set dimension size $d$ to 50, max sequence length $L$ to $200$, and batch size to $128$ in all models. We use an Adam optimizer and tune the learning rate in the range $\{10^{-4},10^{-3},10^{-2}\}$ and set the weight decay to $10^{-5}$. We use the dropout regularization rate of $0.3$ for all models. We set $\gamma=0.5$ in Equation 1, whose reason we will discuss the reason in supplement materials. We define the coarse and fine period to be 10 and 2 days respectively, and we fix the window size to be $m=12$ and $n=4$ for all datasets (§ 4.1.2). We train PrepRec for a maximum of 80 epochs. All experiments are conducted on a Tesla V100 using PyTorch. We repeat each experiment 5 times with different random seeds and report the average performance.
Figure 3. Zero-shot Transfer Results (R@10) with Gaussian noise added to the item popularity statistics (§ 5.3.2). We find that PrepRec is relatively robust to noise.
5.3. Zero-shot Transferability (RQ1)
5.3.1. Zero-shot Transfer Results
We follow the zero-shot inference setting introduced in § 4.3 and report the results in Table 2. We also include the results of PrepRec and the best-performing sequential recommenders trained on the target dataset for reference. In the zero-shot setting, PrepRec shows minimal performance reduction in the target datasets (i.e., 6% maximum and 2% average reduction in R@10). The best zero-shot transfer results from PrepRec only fall short against the selected sequential recommendation baselines by up to 4% and even outperform (by up to 6.5%) them on the Epinions and Office. We found that PrepRec trained on douban-movie and amazon-tools show the highest generalizability, even outperforming the target-trained models on Music (0.811 vs. 0.782 on R@10). We conjecture that this is because Movie is the largest dataset in terms of the number of interactions. Overall, these results show PrepRec ’s effectiveness in zero-shot transfer without any training on interaction data or side information. In addition, this experiment also demonstrates that the popularity dynamics-based item and sequence representations are generalizable across domains.
5.3.2. Robustness to Noise
We further investigate the robustness of PrepRec to possible noise in zero-shot transfer by adding Gaussian noise $\sim\mathcal{N}(0,\sigma)$ to the item popularity statistics and evaluate the zero-shot transfer performance on Douban-Music and Epinions from Douban-Movie. We randomly choose some percentage of items in the sequence to add noise, as indicated in  Figure 3. We find that PrepRec is relatively robust to noise, maintaining robust performance across different noise levels at 20% noised interaction. We attribute this to the model’s ability to learn from the overall popularity dynamics, which is less affected by noise in individual item popularity statistics. In addition, when the noise level is relatively low, e.g., $\sigma\leq 5$, even if 100% of the sequence is noised, PrepRec still holds the performance, indicating significant item popularity shifts exist in the sequence (Figure 1).
@@ 5.3.2. Robustness to Noise
Metric R@10 N@10 R@10 N@10 R@10 N@10
MostPop 0.197 0.139 0.099 0.046 0.163 0.110
SasRec (Kang and McAuley, 2018) 0.749 0.519 0.453 0.291 0.658 0.442
BERT4Rec (Sun et al., 2019) 0.747 0.519 0.461 0.299 0.655 0.456
PrepRec 0.739 0.523 0.443 0.280 0.762 0.551
PrepRec $\Delta$1.3% +0.7% -2.2% -6.3% +15.8% +17.2%
@@
Table 4. Regular sequential recommendation results (§ 5.4.1) with popularity-based negative sampling. PrepRec can learn discriminative item and sequence representations despite depending only on popularity statistics.
5.4. Why Popularity Dynamics? (RQ2)
PrepRec shows excellent performance in zero-shot sequential recommendation. Therefore, we ask: what is the role of popularity dynamics in sequential recommendation, and how much does it explain the performance of state-of-the-art sequential recommenders? Therefore, we propose the following experiments to investigate the importance of popularity dynamics in sequential recommendation.
5.4.1. Regular Sequential Recommendation (RQ2)
We show comparisons of PrepRec against state-of-the-art sequential recommenders in the regular sequential recommendation tasks (Table 3), i.e., all models are trained from scratch. PrepRec achieves competitive performance—within 2% in R@10 and 5% in N@10, with the state-of-the-art baselines. On Epinions, PrepRec even outperforms all baselines by 7.3%, particularly impressive since PrepRec has significantly fewer model parameters (Table 5).
Dataset Office Tool Movie Music Epinions
SasRec 1.331M 3.581M 2.044M 0.542M 1.054M
BERT4Rec 2.687M 7.233M 4.126M 1.094M 2.127M
TiSasRec 1.367M 3.617M 2.127M 0.578M 1.09M
PrepRec 0.045M 0.045M 0.045M 0.045M 0.045M
Table 5. Comparison of model sizes (i.e., number of learnable parameters in millions) over different datasets. PrepRec is 12 to 90x smaller.
PrepRec explicitly models popularity information and the MostPop demonstrates decent performance compared to the remaining baselines, thus we conduct an additional experiment (Table 4) where we sample the unobserved (negative) items based on their popularity (Sun et al., 2019). As shown in Table 4, MostPop’s performance dropped significantly, while PrepRec shows even more competitive performance on some datasets (e.g., Music and Epinions). This suggests PrepRec learns discriminative item and sequence representations.
PrepRec learns item representations through popularity dynamics, which is conceptually different from learning representations specific to each item ID. Therefore, we propose a simple post-hoc interpolation to investigate how much can popularity dynamics explain the performance of state-of-the-art sequential recommenders. We interpolate the scores from PrepRec with the scores from BERT4Rec as follows: $\hat{y}_{intp}(v^{t}_{j}|\mathcal{S}_{u})=\alpha*\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})+(1-\alpha)*\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$, where $\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})$ and $\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ are the scores from PrepRec (Equation 10) and BERT4Rec, respectively. We set $\alpha=0.5$ for all datasets. After interpolation, the performance significantly boosts by up to 34.9% in N@10. Gains are largest in the medium and low-density datasets (Epinions, Amazon), indicating that our model complements existing methods in sparse datasets where item embeddings are less informative. Therefore, it is crucial to consider popularity dynamics to maximize performance.
5.4.2. Qualitative Analysis on Regular Sequential Recommendation
Figure 4. Recommendation results for different item popularity groups (§ 5.4.2), where Group 1 represents the least popular items, and Group 5 represents the most popular items. PrepRec achieves better performance on long-tail items while having competitive performance on popular items.
We analyze the performance of PrepRec in detail. We separate test items into equally sized groups based on their popularity in the training set then compute the average R@10 and N@10 for each group (Figure 4). PrepRec achieves better performance on item group with the least interactions, i.e., long-tail items, while the SasRec and BERT4Rec show stronger performance on popular items. Long-tail item recommendation is a particularly challenging task explored by many previous works (Sankar et al., 2021) and requires recommenders able to learn high-quality representations with just a few interactions. This corresponds to our observation that PrepRec is more robust to data sparsity and can learn discriminative item and sequence representations (§ 5.4.1), showing that long-tail item recommendation can benefit from PrepRec ’s popularity dynamics-based item representations.
5.5. What affects PrepRec performance? (RQ3)
5.5.1. Ablation Study
@@
Metric R@10 N@10 R@10 N@10 R@10 N@10
PrepRec 0.782 0.573 0.536 0.344 0.795 0.580
w/o Relative Time ${\bm{T}}$ (§ 4.1.3) 0.734 0.514 0.541 0.334 0.782 0.562
w/o Positional ${\bm{P}}$ (§ 4.1.4) 0.765 0.544 0.530 0.332 0.772 0.554
w/o Popularity Dynamics $\mathcal{P}$ 0.800 0.594 0.530 0.341 0.761 0.560
w/o Popularity Dynamics $\mathcal{H}$ 0.705 0.582 0.525 0.337 0.730 0.533
Sinuoisal Popularity Encoding 0.779 0.570 0.529 0.340 0.772 0.561
@@
Here, we assess the importance of different components crucial to PrepRec , i.e., relative time encoding (§ 4.1.3), positional encoding (§ 4.1.4), popularity encoder $E_{p}$ (§ 4.1.1), and resolutions for popularity dynamics (§ 4.1.2). We find that removing relative time encoding ${\bm{T}}$ results in the largest performance drop on both the Music and Office datasets. This suggests that the relative time encoding is crucial for effectively capturing the popularity dynamics. Removing positional encoding ${\bm{P}}$ results in a maximum of 2.2% drop in R@10 on the Office dataset, indicating positional encoding is important for capturing sequential information. In addition, changing $E_{p}$ to the non-linear sunusoidal encoding shows worse performance on all datasets, meaning that the linear encoding is more suitable for capturing the popularity dynamics. On the music dataset, removing coarse popularity encoding $\mathcal{P}$ results improves the performance by 2% in R@10, while removing fine popularity encoding $\mathcal{H}$ results in a 7.5% drop in R@10. This suggests that the music domain is more sensitive to recent trends in popularity. Coarse and fine popularity encodings complement each other on other datasets.
@@
Metric R@10 N@10 R@10 N@10 R@10 N@10
$\gamma=0$ (Curr-pop) 0.749 0.542 0.512 0.328 0.689 0.496
$\gamma=0.25$ 0.764 0.529 0.538 0.338 0.761 0.562
$\gamma=0.5^{*}$ (weighted -pop) 0.782 0.573 0.536 0.344 0.795 0.580
$\gamma=0.75$ 0.755 0.520 0.543 0.336 0.747 0.519
$\gamma=1$ (cumul-pop) 0.695 0.452 0.530 0.330 0.733 0.505
Table 7. Recommendation results for varying the discounting factor $\gamma$ in § 4.1.2. $\gamma=0.5$ is the default setting, denoted by ${}^{\prime}*^{\prime}$. We find that $\gamma=0.5$ generally outperforms the other two settings